OpenStudy (anonymous):

5Ln(7x)+4=24 can someone help me with out using Wolfram? lol

OpenStudy (anonymous):

subtract 4, divide by 5 ln(7x)=4 take both sides to base e 7x=e^4 x=(1/7)(e^4)